# Robustness-Inspired Defense Against Backdoor Attacks on Graph Neural Networks

### v1

@@ 4.2 Backdoor Robust GNN Model Training
$\small\min_{f}\mathcal{L}_{f}={\sum}_{v_{i}\in\mathcal{V}_{s}}\log f(v_{i})_{y_{t}}+{\sum}_{v_{j}\in\mathcal{V}_{L}\backslash\mathcal{V}_{s}}\mathcal{L}(f(v_{j}),y_{j}),$ (6)
where $f(v_{i})_{y_{t}}$ denotes the prediction confidence of $f$ for $v_{i}$ on target class $y_{t}$ and $\mathcal{L}(f(v_{j}),y_{j})$ is the cross entropy loss. The resulting classifier is robust against triggers as (i) We explicitly countermeasure the detected backdoors, making our model robust to backdoor. Though there might be target nodes not detected, they generally have smaller prediction variance, meaning that they have less impact on the training of $f$. As these triggers more or less have patterns similar to the detected ones, our training strategy implicitly mitigates their effects. (ii) For triggers that mimic the neighbor distribution of nodes from the target class, our model $f$ is encouraged to explore the subtle differences between clean nodes from the target class and poisoned target nodes, thereby ensuring its clean accuracy. The training algorithm is in Appendix F.
@@ 5.1 Experimental Setup
Datasets. We conduct experiments on three benchmark datasets widely used for node classification, i.e., Cora, Pubmed Sen_Namata_Bilgic_Getoor_Galligher_Eliassi-Rad_2008 , and OGB-arxiv NEURIPS2020_fb60d411 . Cora and Pubmed are small citation networks. OGB-arxiv is a large-scale citation network. More details of the datasets are summarized in Appendix G.1.
Attack Methods. To demonstrate the defense ability of our RIGBD, we evaluate RIGBD on 3 state-of-the-art graph backdoor attack methods, i.e. GTA xi2021graph , UGBA Dai_2023 and DPGBA zhang2024rethinking . The details of these attacks are given in Appendix G.2.
Compared Methods. We implement the backdoor defense strategies Prune Dai_2023 and OD zhang2024rethinking . Additionally, three representative robust GNNs, i.e. RobustGCN zhu2019robust , GNNGuard zhang2020gnnguard and randomized smoothing (RS) wang2021certified with various drop ratios are also selected. We also include ABL li2021anti , which is a popular backdoor defense method in the image domain and aims to train clean models given backdoor-poisoned data. More details of these defense methods are given in Appendix G.3.
Evaluation Protocol. Following existing representative graph backdoor attacks Dai_2023 ; zhang2024rethinking , we conduct experiments on the inductive node classification task. Specifically, we split the graph into two disjoint subgraphs, $\mathcal{G}_{T}$ and $\mathcal{G}_{U}$, with an $80:20$ ratio. The graph $\mathcal{G}_{T}$ is used to train the attacker. Then, the attacker selects target nodes $\mathcal{V}_{B}$ and attaches triggers to these target nodes to form the backdoored graph $\mathcal{G}_{T}$. The number of triggers $|\mathcal{V}_{B}|$ is set as 40, 160, and 565 for Cora, PubMed, and OGB-arxiv, respectively. The trigger size is limited to three nodes in all experiments. The defender trains a model on poisoned graph $\mathcal{G}_{T}$. Next, half of the nodes in $\mathcal{G}_{U}$ are selected as poisoned nodes and are attached with backdoor triggers to test the attack success rate (ASR). The remaining nodes in $\mathcal{G}_{U}$ are kept clean and used to test the clean accuracy (ACC). We also report the Recall and Precision of our method in identifying poisoned nodes. Recall is defined as the percentage of poisoned nodes among the candidate nodes identified, relative to all poisoned nodes. Precision is defined as the percentage of poisoned nodes among the candidate nodes identified. Our RIGBD deploys a 2-layer GCN as the model architecture. Each experiment is conducted 5 times and the average results are reported.
@@
Attacks Defense Cora PubMed OGB-arxiv
GTA GCN 98.98 82.58 93.09 85.18 75.34 65.76
GNNGuard 40.22 78.52 26.93 81.68 0.04 62.58
RobustGCN 90.46 80.37 93.12 81.68 70.95 56.08
Prune 17.63 83.06 28.10 85.05 0.01 63.97
OD 0.04 83.47 0.03 85.27 0.01 65.23
RS 53.14 73.33 42.28 84.58 42.72 58.48
ABL 19.93 81.85 16.18 83.92 11.28 63.24
RIGBD 0.00 83.70 0.01 84.32 0.01 66.51
UGBA GCN 98.76 83.42 96.42 84.64 98.82 63.95
GNNGuard 43.17 78.15 98.97 81.48 95.21 64.61
RobustGCN 98.67 80.00 99.90 82.85 90.35 56.18
Prune 98.89 82.66 92.87 85.09 93.07 62.58
OD 0.03 83.65 0.01 85.19 0.01 65.35
RS 54.24 70.37 44.41 84.68 40.30 58.76
ABL 15.13 81.48 28.60 84.37 34.26 64.93
RIGBD 0.01 84.81 0.01 85.13 0.01 65.21
DPGBA GCN 97.72 83.34 98.63 85.22 95.63 65.72
GNNGuard 85.61 78.52 44.12 80.82 94.66 62.29
RobustGCN 96.68 81.11 94.88 82.54 90.09 60.38
Prune 91.82 85.28 88.64 85.13 90.47 65.53
OD 94.33 83.58 91.32 85.12 93.30 65.47
RS 51.29 69.63 48.83 85.44 41.18 58.44
ABL 86.72 79.26 48.58 79.45 52.56 63.88
RIGBD 0.01 85.19 0.01 84.32 0.00 65.24
@@ 5.2 Performance of Defense
To answer Q1, we compare RIGBD with the baseline defense methods across three datasets. The number of iterations for random edge dropping is set to $K=20$, with a drop ratio of $\beta=0.5$. We report ASR and ACC in Table 2. From the table, we observe: (i) Across all datasets and attack methods, RIGBD consistently achieves the lowest ASR score, typically close to $0\%$. Although defense methods such as Prune and OD perform well against GTA and UGBA attacks, they fail to defend against DPGBA, which generates in-distribution triggers. This indicates that RIGBD is highly effective in defending against various types of backdoor triggers and backdoor attacks. (ii) Our RIGBD achieves comparable or slightly better clean accuracy compared to the vanilla GCN. This indicates that our approach of random edge dropping typically leads to higher prediction variance for poisoned nodes, effectively identifying backdoor triggers. In cooperation with this precise identification of poisoned nodes, our strategy to train a backdoor robust model using Eq. (6) can significantly reduce the ASR while maintaining clean accuracy. Additional results on RS with varying drop ratios and RIGBD using different GNN architectures can be found in Appendix I.
@@ 5.3 Ability to Detect Poisoned Nodes
To answer Q2, we present the recall and precision of RIGBD in identifying poisoned nodes. Using the same setting as in Sec. 5.2, the results on OGB-arxiv are shown in Table 3. We also include the corresponding ASR and ACC to illustrate the correlation between detection ability and defense performance. Clean accuracy (Clean ACC), obtained from a model trained on the clean graph, is provided as a reference to evaluate the model’s performance on clean nodes. From the table, we observe the following: (i) Our RIGBD demonstrates consistently high precision, always over 90%, in detecting poisoned nodes across three attack methods, with detection recall always exceeding 80%. This indicates that our strategy of random edge dropping typically leads to higher prediction variance for poisoned nodes. (ii) Although the recall for the GTA attack method is 84.9%, we still achieve an ASR close to 0% while maintaining ACC. This verifies the stability of our strategy to train a robust node classifier using Eq. (6). Even when some poisoned nodes are not identified, it still efficiently helps the model unlearn the triggers. More results on other datasets are in Appendix J.
@@ 5.4 Hyperparameter Analysis and Ablation Study
Hyperparameter Analysis. To answer Q3, we conduct experiments to show how different drop ratios $\beta$ and different numbers of iterations of random edge dropping $K$ impact the performance of RIGBD. Specifically, we vary the value of $K$ as $\left\{2,5,25,50,100\right\}$, and the values of $\beta$ as $\left\{0.1,0.3,0.5,0.7,0.9\right\}$. The attack method used is DPGBA. The other settings are the same as the evaluation protocol in Sec. 5.1. The results on OGB-arxiv are shown in Fig. 3. From the figure, we observe: (i) Our RIGBD is stable in terms of ASR and Precision. Notably, when the drop ratio $\beta=0.1$ and $K=2$, the recall of poisoned node detection is around 17.4%. However, we still achieve robust defense performance with 1.86% ASR. This is because our framework identifies the most efficient triggers that lead to large prediction variance and focuses on minimizing their impact. By doing so, the influence of less efficient triggers is inherently counteracted. In contrast, when we simply remove those 17.4% triggers from the dataset and train a normal GNN on the remaining dataset, the ASR is around 80%. (ii) As $K$ increases, the recall also increases, empirically verifying our theorems that the clean nodes will have stable node embeddings and predictions. As $\beta$ increases from 0.1 to 0.7, the recall also increases without a decrease in precision. When $\beta=0.9$, only about 10% of edges remain in each iteration. Although the recall decreases slightly, we still achieve consistently high precision. This further shows the robustness of RIGBD against various chosen hyperparameters. More results on the hyperparameter analysis for ACC are in Appendix K.
@@
Performance Under Different Number of Triggers. The performance and analysis of RIGBD against different attack budgets is provided in the Appendix M.
@@ A.1 Graph Backdoor Attacks
Backdoor attacks have been extensively studied in the image domain li2022backdoor ; gu2019badnets ; chen2017targeted ; sun2023efficient ; li2023embarrassingly . Initial works focused on directly poisoning training samples liu2020reflection ; chen2017targeted , while others explored making the triggers invisible li2021invisible ; NEURIPS2021_9d99197e . Additionally, hidden backdoors can be embedded through transfer learning kurita2020weight , modifying model parameters chen2021proflip , and adding extra malicious modules tang2020embarrassingly . Recently, studies have begun investigating backdoor attacks on GNNs, which differ from the more common poisoning and evasion attacks. Backdoor attacks involve injecting malicious triggers into the training data, causing the model to make incorrect predictions when these triggers are present in test samples. This type of attack subtly manipulates the training phase, ensuring the model performs as expected under normal conditions but fails when trigger-embedded inputs are present. Among pioneering efforts, SBA zhang2021backdoor introduced a method for injecting universal triggers into training samples using a subgraph-based approach, though its attack success rate was low. GTA xi2021graph advanced this by developing a technique for generating adaptive triggers, customizing perturbations for individual samples to enhance attack effectiveness. In UGBA Dai_2023 , an algorithm for selecting poisoned nodes was introduced to optimize the attack budget, along with an adaptive trigger generator to create triggers with high cosine similarity to the target node. DPGBA zhang2024rethinking demonstrate that existing backdoor attack methods on graphs suffer from either a low attack success rate or outlier issues. To address these problems, they propose an adversarial learning strategy to generate in-distribution triggers and introduce a novel loss function to enhance the attack success rate with these in-distribution triggers.
@@ A.2 Graph Backdoor Defense
Backdoor defenses have been extensively investigated in the context of non-structured data. To address the threat of backdoor attacks, various defenses have been proposed, which can be broadly classified into two main categories: empirical backdoor defenses and certified backdoor defenses. Empirical backdoor defenses wang2019neural ; kolouri2020universal ; li2021anti are based on specific observations or understandings of existing attacks and generally perform well in practice; however, they lack theoretical guarantees and may be circumvented by adaptive attacks. On the other hand, certified backdoor defenses wang2020certifying ; weber2023rab ; xie2021crfl offer theoretical guarantees of their validity under certain assumptions, but their practical performance is often weaker than empirical defenses due to the difficulty in meeting these assumptions. Improving defenses against backdoor attacks remains a significant open challenge. Limited defenses have been proposed against the graph backdoor. UGBA Dai_2023 highlights that in attack methods GTA xi2021graph , the attributes of triggers differ significantly from the attached poisoned nodes, violating the homophily property typically observed in real-world graphs. To address this, they propose a defense method called Prune, which involves removing edges that connect nodes with low cosine similarity. This approach significantly reduces ASR of previous works. DPGBA zhang2024rethinking further notes that although the generated triggers in UGBA Dai_2023 may demonstrate high cosine similarity to target nodes, the triggers in both UGBA Dai_2023 and GTA xi2021graph are still outliers. Thus, they propose a defense method called OD, which involves training a graph auto-encoder and filtering out nodes with high reconstruction loss. Their empirical results show that existing backdoor attack methods on graphs suffer from either a low ASR or outlier issues. However, there is no existing work on defending against DPGBA, which generates in-distribution triggers.

### v2

@@ 4.2 Backdoor robust GNN model training
$\min_{f}\mathcal{L}_{f}={\sum}_{v_{i}\in\mathcal{V}_{s}}\log f(v_{i})_{y_{t}}+{\sum}_{v_{j}\in\mathcal{V}_{L}\backslash\mathcal{V}_{s}}\mathcal{L}(f(v_{j}),y_{j}),$ (6)
where $f(v_{i})_{y_{t}}$ denotes the prediction confidence of $f$ for $v_{i}$ on target class $y_{t}$ and $\mathcal{L}(f(v_{j}),y_{j})$ is the cross entropy loss. The resulting classifier is robust against triggers as (i) We explicitly countermeasure the detected backdoors, making our model robust to backdoor. Though there might be target nodes not detected, they generally have smaller prediction variance, meaning that they have less impact on the training of $f$. As these triggers more or less have patterns similar to the detected ones, our training strategy implicitly mitigates their effects. (ii) For triggers that mimic the neighbor distribution of nodes from the target class, our model $f$ is encouraged to explore the subtle differences between clean nodes from the target class and poisoned target nodes, thereby ensuring its clean accuracy. The training algorithm is in Appendix G.
@@ 5.1 Experimental setup
Datasets. We conduct experiments on three benchmark datasets widely used for node classification, i.e., Cora, Citeseer, Pubmed (Sen et al., 2008), Physics (Sinha et al., 2015), Flickr (Zeng et al., 2019) and OGB-arxiv (Hu et al., 2020). More details of the datasets are summarized in Appendix H.1.
Attack Methods. To demonstrate the defense ability of our RIGBD, we evaluate RIGBD on 3 state-of-the-art graph backdoor attack methods, i.e. GTA (Xi et al., 2021), UGBA (Dai et al., 2023) and DPGBA (Zhang et al., 2024). The details of these attacks are given in Appendix H.2.
Compared Methods. We implement the backdoor defense strategies Prune (Dai et al., 2023) and OD (Zhang et al., 2024). Additionally, three representative robust GNNs, i.e. RobustGCN (Zhu et al., 2019), GNNGuard (Zhang & Zitnik, 2020) and randomized smoothing (RS) (Wang et al., 2021) with various drop ratios are also selected. We also include ABL (Li et al., 2021a), which is a popular backdoor defense method in the image domain and aims to train clean models given backdoor-poisoned data. More details of these defense methods are given in Appendix H.3.
Evaluation Protocol. Following existing representative graph backdoor attacks (Dai et al., 2023; Zhang et al., 2024), we split the graph into two disjoint subgraphs, $\mathcal{G}_{T}$ and $\mathcal{G}_{U}$, with an $80:20$ ratio. The graph $\mathcal{G}_{T}$ is used to train the attacker. Then, the attacker selects target nodes $\mathcal{V}_{B}$ and attaches triggers to these target nodes to form the backdoored graph $\mathcal{G}_{T}$. The number of triggers $|\mathcal{V}_{B}|$ is set as 40, 80, 160, 160, 160 and 565 for Cora, Citeseer, PubMed, Physics, Flickr, and OGB-arxiv, respectively. The trigger size is limited to three nodes in all experiments. The defender trains a model on poisoned graph $\mathcal{G}_{T}$. Next, half of the nodes in $\mathcal{G}_{U}$ are selected as poisoned nodes and are attached with backdoor triggers to test the attack success rate (ASR). The remaining nodes in $\mathcal{G}_{U}$ are kept clean and used to test the clean accuracy (ACC). We also report the Recall and Precision of our method in identifying poisoned nodes. Recall is defined as the percentage of poisoned nodes among the candidate nodes identified, relative to all poisoned nodes. Precision is defined as the percentage of poisoned nodes among the candidate nodes identified. Our RIGBD deploys a 2-layer GCN as the model architecture. Each experiment is conducted 5 times and the average results are reported.
@@
Attacks Defense Cora Citeseer PubMed Physics Flickr OGB-arxiv
ASR(%) $\downarrow$ ACC(%) $\uparrow$ ASR(%) $\downarrow$ ACC(%) $\uparrow$ ASR(%) $\downarrow$ ACC(%) $\uparrow$ ASR(%) $\downarrow$ ACC(%) $\uparrow$ ASR(%) $\downarrow$ ACC(%) $\uparrow$ ASR(%) $\downarrow$ ACC(%) $\uparrow$
GTA GCN 98.98 82.58 100.00 73.70 93.09 85.18 96.42 89.30 88.45 41.23 75.34 65.76
GNNGuard 40.22 78.52 55.26 63.55 26.93 81.68 43.87 78.72 0.00 40.40 0.04 62.58
RobustGCN 90.46 80.37 95.31 73.79 93.12 81.68 95.47 94.84 85.42 42.26 70.95 56.08
Prune 17.63 83.06 12.24 72.46 28.10 85.05 8.34 88.45 12.56 41.27 0.01 63.97
OD 0.04 83.47 0.04 72.84 0.03 85.27 0.12 90.24 2.12 41.42 0.01 65.23
RS 53.14 73.33 52.86 65.66 42.28 84.58 57.70 92.52 52.85 42.31 42.72 58.48
ABL 19.93 81.85 13.09 73.19 16.18 83.92 16.82 92.06 17.54 41.46 11.28 63.24
RIGBD 0.00 83.70 0.34 74.10 0.01 84.32 0.32 95.10 0.00 44.21 0.01 66.51
UGBA GCN 98.76 83.42 100.00 74.70 96.42 84.64 100.00 95.94 93.14 44.71 98.82 63.95
GNNGuard 43.17 78.15 94.53 64.76 98.97 81.48 95.26 87.54 96.93 42.15 95.21 64.61
RobustGCN 98.67 80.00 99.82 71.69 99.90 82.85 99.94 94.08 32.17 41.82 90.35 56.18
Prune 98.89 82.66 97.68 74.35 92.87 85.09 94.67 93.87 91.43 43.65 93.07 62.58
OD 0.03 83.65 0.06 73.80 0.01 85.19 0.02 95.36 1.65 43.57 0.01 65.35
RS 54.24 70.37 50.34 69.88 44.41 84.68 47.77 94.29 20.69 42.18 40.30 58.76
ABL 15.13 81.48 12.08 73.19 28.60 84.37 14.87 94.69 15.32 41.66 34.26 64.93
RIGBD 0.01 84.81 0.00 73.80 0.01 85.13 0.12 95.71 0.00 43.66 0.01 65.21
DPGBA GCN 97.72 83.34 100.00 74.09 98.63 85.22 100.00 95.59 90.79 44.87 95.63 65.72
GNNGuard 85.61 78.52 46.98 60.84 44.12 80.82 88.72 88.76 95.85 43.52 94.66 62.29
RobustGCN 96.68 81.11 91.64 71.08 94.88 82.54 91.25 94.87 24.60 41.69 90.09 60.38
Prune 91.82 85.28 94.80 73.21 88.64 85.13 94.27 94.73 88.96 44.75 90.47 65.53
OD 94.33 83.58 98.42 73.66 91.32 85.12 98.72 95.48 90.42 43.63 93.30 65.47
RS 51.29 69.63 50.34 71.08 48.83 85.44 48.19 95.30 27.94 42.21 41.18 58.44
ABL 86.72 79.26 11.41 73.49 48.58 79.45 11.74 95.30 30.42 41.71 52.56 63.88
RIGBD 0.01 85.19 0.33 73.79 0.03 84.56 0.21 95.79 0.00 43.78 0.00 65.24
@@ 5.2 Performance of defense
To answer Q1, we compare RIGBD with the baseline defense methods across three datasets. The number of iterations for random edge dropping is set to $K=20$, with a drop ratio of $\beta=0.5$. We report ASR and ACC in Table 2. From the table, we observe: (i) Across all datasets and attack methods, RIGBD consistently achieves the lowest ASR score, typically close to $0\%$. Although defense methods such as Prune and OD perform well against GTA and UGBA attacks, they fail to defend against DPGBA, which generates in-distribution triggers. This indicates that RIGBD is highly effective in defending against various types of backdoor triggers and backdoor attacks. (ii) Our RIGBD achieves comparable or slightly better clean accuracy compared to the vanilla GCN. This indicates that our approach of random edge dropping typically leads to higher prediction variance for poisoned nodes, effectively identifying backdoor triggers. In cooperation with this precise identification of poisoned nodes, our strategy to train a backdoor robust model using Eq. 6 can significantly reduce the ASR while maintaining clean accuracy. Additional results on RS with varying drop ratios and RIGBD using different GNN architectures can be found in Appendix J.
@@ 5.3 Ability to detect poisoned nodes
To answer Q2, we present the recall and precision of RIGBD in identifying poisoned nodes. Using the same setting as in Sec. 5.2, the results on OGB-arxiv are shown in Table 3. We also include the corresponding ASR and ACC to illustrate the correlation between detection ability and defense performance. Clean accuracy (Clean ACC), obtained from a model trained on the clean graph, is provided as a reference to evaluate the model’s performance on clean nodes. From the table, we observe the following: (i) Our RIGBD demonstrates consistently high precision, always over 90%, in detecting poisoned nodes across three attack methods, with detection recall always exceeding 80%. This indicates that our strategy of random edge dropping typically leads to higher prediction variance for poisoned nodes. (ii) Although the recall for the GTA attack method is 84.9%, we still achieve an ASR close to 0% while maintaining ACC. This verifies the stability of our strategy to train a robust node classifier using Eq. 6. Even when some poisoned nodes are not identified, it still efficiently helps the model unlearn the triggers. More results on other datasets are in Appendix L.
@@ 5.4 Hyperparameter Analysis and Ablation Study
Hyperparameter Analysis. To answer Q3, we conduct experiments to show how different drop ratios $\beta$ and different numbers of iterations of random edge dropping $K$ impact the performance of RIGBD. Specifically, we vary the value of $K$ as $\left\{2,5,25,50,100\right\}$, and the values of $\beta$ as $\left\{0.1,0.3,0.5,0.7,0.9\right\}$. The attack method used is DPGBA. The other settings are the same as the evaluation protocol in Sec. 5.1. The results on OGB-arxiv are shown in Fig. 3. From the figure, we observe: (i) Our RIGBD is stable in terms of ASR and Precision. Notably, when the drop ratio $\beta=0.1$ and $K=2$, the recall of poisoned node detection is around 17.4%. However, we still achieve robust defense performance with 1.86% ASR. This is because our framework identifies the most efficient triggers that lead to large prediction variance and focuses on minimizing their impact. By doing so, the influence of less efficient triggers is inherently counteracted. In contrast, when we simply remove those 17.4% triggers from the dataset and train a normal GNN on the remaining dataset, the ASR is around 80%. (ii) As $K$ increases, the recall also increases, empirically verifying our theorems that the clean nodes will have stable node embeddings and predictions. As $\beta$ increases from 0.1 to 0.7, the recall also increases without a decrease in precision. When $\beta=0.9$, only about 10% of edges remain in each iteration. Although the recall decreases slightly, we still achieve consistently high precision. This further shows the robustness of RIGBD against various chosen hyperparameters. More results on the hyperparameter analysis for ACC are in Appendix M.
@@
Impact of Various Numbers of Triggers and Trigger Sizes. The performance and analysis of RIGBD against various attack budgets and trigger sizes are provided in Appendix O and P.
@@ 6 Conclusion
Acknowledgment
This material is based upon work supported by, or in part by the Army Research Office (ARO) under grant number W911NF-21-10198, the Department of Homeland Security (DHS) under grant number 17STCIN00001-05-00, and Cisco Faculty Research Award. The views and conclusions contained in this material are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of the funding agencies.
@@ A.1 Graph backdoor attacks
Backdoor attacks have been extensively studied in the image domain (Li et al., 2022; Gu et al., 2019; Chen et al., 2017; Sun et al., 2023; Li et al., 2023). Initial works focused on directly poisoning training samples (Liu et al., 2020; Chen et al., 2017; Xing et al., 2024), while others explored making the triggers invisible (Li et al., 2021b; Doan et al., 2021). Additionally, hidden backdoors can be embedded through transfer learning (Kurita et al., 2020), modifying model parameters (Chen et al., 2021), and adding extra malicious modules (Tang et al., 2020). Recently, studies have begun investigating backdoor attacks on GNNs (Dai et al., 2023; Xi et al., 2021; Zhang et al., 2021; 2024; Lin et al., 2024), which differ from the more common poisoning and evasion attacks (Dai et al., 2024; 2022a; Ma et al., 2021b). Backdoor attacks involve injecting malicious triggers into the training data, causing the model to make incorrect predictions when these triggers are present in test samples. This type of attack subtly manipulates the training phase, ensuring the model performs as expected under normal conditions but fails when trigger-embedded inputs are present. Among pioneering efforts, SBA (Zhang et al., 2021) introduced a method for injecting universal triggers into training samples using a subgraph-based approach, though its attack success rate was low. GTA (Xi et al., 2021) advanced this by developing a technique for generating adaptive triggers, customizing perturbations for individual samples to enhance attack effectiveness. In UGBA (Dai et al., 2023), an algorithm for selecting poisoned nodes was introduced to optimize the attack budget, along with an adaptive trigger generator to create triggers with high cosine similarity to the target node. DPGBA (Zhang et al., 2024) demonstrate that existing backdoor attack methods on graphs suffer from either a low attack success rate or outlier issues. To address these problems, they propose an adversarial learning strategy to generate in-distribution triggers and introduce a novel loss function to enhance the attack success rate with these in-distribution triggers. In this paper, we investigate defense mechanisms against backdoor attacks that involve attaching backdoor triggers to target nodes (Xi et al., 2021; Dai et al., 2023; Zhang et al., 2024). These attacks are considered a substantial threat to real-world applications of GNNs, especially in comparison to attacks involving direct manipulation of node attributes (Xing et al., 2024). In real-world contexts, such as social media networks, it is more practical for an adversary to introduce malicious nodes (e.g., fake accounts) and connect them to a target node rather than altering the target node’s attributes. Thus, it is anticipated that this form of attack will receive increasing research attention. Given this threat, it is crucial to develop robust defense frameworks capable of mitigating these types of attacks.
@@ A.2 Graph backdoor defense
Backdoor defenses have been extensively investigated in the context of non-structured data. To address the threat of backdoor attacks, various defenses have been proposed, which can be broadly classified into two main categories: empirical backdoor defenses and certified backdoor defenses. Empirical backdoor defenses (Wang et al., 2019; Kolouri et al., 2020; Li et al., 2021a) are based on specific observations or understandings of existing attacks and generally perform well in practice; however, they lack theoretical guarantees and may be circumvented by adaptive attacks. On the other hand, certified backdoor defenses (Wang et al., 2020; Weber et al., 2023; Xie et al., 2021) offer theoretical guarantees of their validity under certain assumptions, but their practical performance is often weaker than empirical defenses due to the difficulty in meeting these assumptions. Improving defenses against backdoor attacks remains a significant open challenge. Limited defenses have been proposed against the graph backdoor. UGBA (Dai et al., 2023) highlights that in attack methods GTA (Xi et al., 2021), the attributes of triggers differ significantly from the attached poisoned nodes, violating the homophily property typically observed in real-world graphs. To address this, they propose a defense method called Prune, which involves removing edges that connect nodes with low cosine similarity. This approach significantly reduces ASR of previous works. DPGBA (Zhang et al., 2024) further notes that although the generated triggers in UGBA (Dai et al., 2023) may demonstrate high cosine similarity to target nodes, the triggers in both UGBA (Dai et al., 2023) and GTA (Xi et al., 2021) are still outliers. Thus, they propose a defense method called OD, which involves training a graph auto-encoder and filtering out nodes with high reconstruction loss. Their empirical results show that existing backdoor attack methods on graphs suffer from either a low ASR or outlier issues. However, there is no existing work on defending against DPGBA, which generates in-distribution triggers.